OpenStudy (anonymous):

If n=2, and l=0, then what are the possible values of ml ?

OpenStudy (anonymous):

n =2, then l can be 0,1 and 2. since l=0, then, ml= 2l+1 = 2*0+1=1 this makes sense as you think of it, l=0 is a s-orbital, which is spherical shaped and have no orientation preferences.